OpenStudy (anonymous):

1. Are the complex fractions below equivalent? 2. What is the excluded value of the rational expression

OpenStudy (anonymous):

Here is the term for 1.

OpenStudy (anonymous):

Here is the term for 2.

OpenStudy (anonymous):

Hate math....

OpenStudy (anonymous):

me 2

OpenStudy (anonymous):

@wackett1996 please help me!

OpenStudy (anonymous):

@hartnn please help me!

hartnn (hartnn):

When you divide by some thing, you multiply by its reciprocal., \(\huge \dfrac{♥}{☺}=♥\times \dfrac{1}{☺}\) or in more general, \(\Huge \dfrac{\dfrac{☻}{♣}}{\dfrac{♠}{♥}}=\dfrac{☻}{♣}\times \dfrac{♥}{♠}\) notice how ♥ came to numerator and ♠ became the denominator. so, for 1)  and  Are they same ??? Applying same thing for 2) \(\huge \dfrac{x-4}{6} \div \dfrac{x+1}{7} = \dfrac{x-4}{6} \times \dfrac{7}{x+1}=...?\)